Susan invested part of her $24,000 bonus in a fund that paid 9% profit and invested the rest in stock that suffered a 5% loss. Find the amount of each investment if her overall net profit was $480.
I can't figure out the formula or how to set this problem up.
Thank you

Answer by checkley77(12844)

RATE1*AMOUNT1+RATE2*AMOUNT2=TOTAL INTEREST
------------------------------------------
.09x-.05(24,000-x)=480
.09x-1,200+.05x=480
.09x+.05x=480+1,200
.14x=1,680
x=1,680/.14
x=12,000 amount invested @ 9%.
24,000-12,000=12,000 amount invested @ 4%.
Proof:
.09*12,000-.05*12,000=480
1080-600=480
480=480